# OptiState: State Estimation of Legged Robots using Gated Networks with Transformer-based Vision and Kalman Filtering

### v2

@@ II-A Problem Definition
$\mathbf{O}_{k}=\delta($ $\hat{\mathbf{x}}_{k-N\mathrel{\mathop{:}}k},\mathbf{L}_{k-N\mathrel{\mathop{:}}k}^{\rm depth},\mathbf{p}_{k-N\mathrel{\mathop{:}}k},\dot{\mathbf{p}}_{k-N\mathrel{\mathop{:}}k},$ (2)
$\mathbf{a}^{\rm imu}_{k-N\mathrel{\mathop{:}}k},\boldsymbol{\alpha}^{\rm imu}_{k-N\mathrel{\mathop{:}}k},\mathbf{f}_{k-N\mathrel{\mathop{:}}k})$
where $\delta$ is a function describing the GRU (see Sec. II-E), with the following inputs: $\hat{\mathbf{x}}_{k-N\mathrel{\mathop{:}}k}$, which is the estimated trunk state from (1), $\mathbf{L}_{k-N\mathrel{\mathop{:}}k}^{\rm depth}$ is the latent space of the depth image from our ViT described in Sec. II-E, $\mathbf{p}_{k-N\mathrel{\mathop{:}}k}$ and $\dot{\mathbf{p}}_{k-N\mathrel{\mathop{:}}k}$ are the estimated footstep positions and velocities (obtained from the joint encoder positions and velocities, as explained in Sec. II-B), $\mathbf{a}^{\rm imu}$ and $\boldsymbol{\alpha}^{\rm imu}$ are the IMU linear and angular accelerations respectively (we use these IMU accelerations because they may indicate instances of slipping or sudden falling), $\mathbf{f}_{k-N\mathrel{\mathop{:}}k}$ are the same ground reaction forces described in (1), the GRU output is given by $\mathbf{O}_{k}=[\bar{\mathbf{x}}_{k},\boldsymbol{\mu}_{k}]$, where $\bar{\mathbf{x}}_{k}$ is the updated trunk state and $\boldsymbol{\mu}_{k}$ is the absolute error between the GRU prediction and ground truth ($\mathbf{x}^{\rm mocap}$) written as $|\bar{\mathbf{x}}_{k}-\mathbf{x}^{\rm mocap}_{k}|$, and we use $N$ time steps to indicate the recurrent nature of the GRU. We include ground reaction forces as inputs as they contain information on the reference trajectory and contact sequence from the MPC optimization, and depth images to convey robot height and environmental semantics in our training procedure.
@@ II-B Measurements
$\mathbf{v}^{\rm odom}=-\frac{1}{n_{c}}\sum_{i=1}^{n}\mathbf{R}_{b}^{w}(\dot{\mathbf{p}}^{i}-\boldsymbol{\omega}^{\rm imu}\times\mathbf{p}^{i})\text{ (foot $i$ in contact)}$) (3)
@@ II-C Model
$\mathbf{Dyn}$ $(\mathbf{x}_{k},\mathbf{f}_{k})=\begin{bmatrix}\boldsymbol{\theta}_{k}+\Delta T\mathbf{R}_{b}^{w\top}\boldsymbol{\omega}_{k}\\ \mathbf{r}_{{\rm}k}+\Delta T\mathbf{v}_{{\rm}k}\\ \boldsymbol{\omega}_{k}+\Delta T(\sum_{i=1}^{n}\hat{\mathbf{I}}^{-1}\left[\mathbf{p}^{b}_{i,k}\right]_{\times}\mathbf{f}^{i}_{k})\\ \mathbf{v}_{k}+\Delta T(\sum_{i=1}^{n}\frac{\mathbf{f}^{i}_{k}}{m}+\mathbf{g})\end{bmatrix}$ (5f)
where $\mathbf{x}_{k}$ is the trunk state of the robot, the ground reaction forces are represented by $\mathbf{f}_{k}^{i}$ for each foot $i$, $m$ is the total mass of the robot, $\mathbf{g}$ is the gravity vector, $\mathbf{\hat{I}}$ is the inertia tensor matrix in world coordinates, $\mathbf{p}_{i,k}^{b}$ is the footstep position of foot $i$ in the body frame, $\Delta T$ is the discretized time step, and lastly, $[\bullet]_{\times}$ is defined as the skew-symmetric matrix (same definition as used in [6]). We can rewrite (II-C):
@@
During operation of the robot, the MPC QP outputs ground reaction forces of the feet in contact with the ground, which are assigned into joint motor torques. These outputs enable the robot trunk to follow a reference state over a time horizon [6]. Thus, we can simply use these forces with the trunk state and footstep positions at time step $k$ in the model described by (II-C), to predict the trunk state at time step $k+1$.

### v3

@@ II-A Problem Definition
$\mathbf{O}_{k}=\delta($ $\hat{\mathbf{x}}_{k-N:k},\mathbf{L}_{k-N:k}^{\rm depth},\mathbf{p}_{k-N:k},\dot{\mathbf{p}}_{k-N:k},$ (2)
$\mathbf{a}^{\rm imu}_{k-N:k},\boldsymbol{\alpha}^{\rm imu}_{k-N:k},\mathbf{f}_{k-N:k})$
where $\delta$ is a function describing the GRU (see Sec. II-E), with the following inputs: $\hat{\mathbf{x}}_{k-N:k}$, which is the estimated trunk state from (1), $\mathbf{L}_{k-N:k}^{\rm depth}$ is the latent space of the depth image from our ViT described in Sec. II-E, $\mathbf{p}_{k-N:k}$ and $\dot{\mathbf{p}}_{k-N:k}$ are the estimated footstep positions and velocities (obtained from the joint encoder positions and velocities, as explained in Sec. II-B), $\mathbf{a}^{\rm imu}$ and $\boldsymbol{\alpha}^{\rm imu}$ are the IMU linear and angular accelerations respectively (we use these IMU accelerations because they may indicate instances of slipping or sudden falling), $\mathbf{f}_{k-N:k}$ are the same ground reaction forces described in (1), the GRU output is given by $\mathbf{O}_{k}=[\bar{\mathbf{x}}_{k},\boldsymbol{\mu}_{k}]$, where $\bar{\mathbf{x}}_{k}$ is the updated trunk state and $\boldsymbol{\mu}_{k}$ is the absolute error between the GRU prediction and ground truth ($\mathbf{x}^{\rm mocap}$) written as $|\bar{\mathbf{x}}_{k}-\mathbf{x}^{\rm mocap}_{k}|$, and we use $N$ time steps to indicate the recurrent nature of the GRU. We include ground reaction forces as inputs as they contain information on the reference trajectory and contact sequence from the MPC optimization, and depth images to convey robot height and environmental semantics in our training procedure.
@@ II-B Measurements
$\mathbf{v}^{\rm odom}=-\frac{1}{n_{c}}\sum_{i=1}^{n}\mathbf{R}_{b}^{w}(\dot{\mathbf{p}}^{i}+\boldsymbol{\omega}^{\rm imu}\times\mathbf{p}^{i})\text{ (foot $i$ in contact)}$) (3)
@@ II-C Model
$\mathbf{Dyn}$ $(\mathbf{x}_{k},\mathbf{f}_{k})=\begin{bmatrix}\boldsymbol{\theta}_{k}+\Delta T\mathbf{R}_{b}^{w\top}\boldsymbol{\omega}_{k}\\ \mathbf{r}_{{\rm}k}+\Delta T\mathbf{v}_{{\rm}k}\\ \boldsymbol{\omega}_{k}+\Delta T(\sum_{i=1}^{n}\hat{\mathbf{I}}^{-1}\left[\mathbf{p}^{b}_{i,k}\right]_{\times}\mathbf{f}^{i}_{k})\\ \mathbf{v}_{k}+\Delta T(\sum_{i=1}^{n}\frac{\mathbf{f}^{i}_{k}}{m}+\mathbf{g})\end{bmatrix}$ (5b)
where $\mathbf{x}_{k}$ is the trunk state of the robot, the ground reaction forces are represented by $\mathbf{f}_{k}^{i}$ for each foot $i$, $m$ is the total mass of the robot, $\mathbf{g}$ is the gravity vector, $\mathbf{\hat{I}}$ is the inertia tensor matrix in world coordinates, $\mathbf{p}_{i,k}^{b}$ is the footstep position of foot $i$ in the body frame, $\Delta T$ is the discretized time step, and lastly, $[\bullet]_{\times}$ is defined as the skew-symmetric matrix (same definition as used in [6]). We can rewrite (5):
@@
During operation of the robot, the MPC QP outputs ground reaction forces of the feet in contact with the ground, which are assigned into joint motor torques. These outputs enable the robot trunk to follow a reference state over a time horizon [6]. Thus, we can simply use these forces with the trunk state and footstep positions at time step $k$ in the model described by (5), to predict the trunk state at time step $k+1$.